Saturday 03/17/12

11:00 A – 10:00P

  • Experimenting with heat fixing embryos
  • working on ACS application
  • filter extract new tracer beads.  Initial test suggests still contains raw dyes.  Tried spin.
  • No obvious colonies on plates.  Back at 37C O/N.
  • STORM imaging of Pho K27
This entry was posted in Summaries and tagged , . Bookmark the permalink.